AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 185     Entry time: Thu Feb 3 22:05:33 2022
Author: JM, TD, OH(remotely) 
Subject: Bias update and grounding 

9:30 Bias remained stable overnight - 60V 2.95 uA.

10:00 No grounding attached initailly. Pulser ON, lots of noise across all ~250-300 kHz across biased detector. Waveforms and histograms saved to /tmp. Pulser width for aida 6 ~300 channels. Aida 7,8 similar, aida 5 producing no ADC data. 

11:00 Grounding connected as in elog 181. Rates decreased to ~180-250 kHz. Waveforms and histograms saved to /tmp. 

12:00 Bias slowly increased too 100V giving a leakage current of ~3.5 uA which was stable for several hours. Bias varied between 60V and 100V for noise reduction perposes, 80V deemed least noisy at the time, noise still ~180-200 kHz

13:30 New grounding trialled. Bottom detector removed from pulser and grounding loops so that only the biased detector is grounded. Bias set to 80V

14:00 Rates significantly reduced for aida 6,7,8 - aida5 still producing no ADC data. Rates are in range 15-50 kHZ with noisy channels accounting for most of the rates. Waveforms and histograms saved to /tmp. Pulser peak on aida 6 ~300 channels. 

16:00 Aida 5 fee investigated to reason it is not producing data. FEE was not properly connected to ERNI connector due to incorrect installation of black adaptor card mount. Adaptor card mount reinstalled and fee remounted. Aida 5 now produces ADC with similar rates to aida6,7,8. 

During investigation of aida5, aida4 and 5 were swapped. Config files relating to MAC addresses, polarity etc of FEES has been updated. Elog 175 will be appended to include this swap.  

19:00 Bias set to 80V, DAQ running with pulser off to allow for alpha detection overnight. 

ELOG V3.1.4-unknown